Mr. Smith purchased $125.00 worth of spices and herbs. Because this was such a large order, the supplier charged Mr. Smith only $95.00. What percent discount did Mr. smith recieve?
Answer by stanbon(48535)

Mr. Smith purchased $125.00 worth of spices and herbs. Because this was such a large order, the supplier charged Mr. Smith only $95.00. What percent discount did Mr. smith recieve?
---------------------------------
Value of discount = 125-95 = 30
----
Percent of discount 30/125 = 24%
========================================
Cheers,
Stan H.